OpenStudy (anonymous):

How much impulse is needed to stop a ball that is 1.0 kg in mass and 9.0 m/s in speed? Please explain how you figure it out so I can do other ones similar to this.

OpenStudy (goformit100):

For Solving this Question, I would like to ask you that, do you have the detailed conception of : impulse

OpenStudy (anonymous):

stop==> final velocity =0ms-1 so, by \(F=\frac{m(v-u)}{t}\), where impulse=Ft=m(v-u), sub everything in. do you need further help?